# Direct Quantized Training of Language Models with Stochastic Rounding

### v1

@@ 4.5 Impact of the Bit Width in DQT
Figure 3: Comparison of different bit widths in DQT. Higher $n$-bit results in better performance. Optimal learning rates are selected for each model: 1e-3 for 1.58-bit and 3-bit models, and 5e-4 for 4-bit and 8-bit models.
@@
Next, we proceed to examine the impact of varying the number of bits in DQT. Specifically, we experiment with $n$ in {1.58, 3, 4, 8}, and the corresponding results are presented in Figure 3. We select the optimal learning rate for each model.
@@
From Figure 3, we can first observe a clear trend: as the number of bits used in DQT increases, the model’s performance improves consistently. Notably, for our 1.58-bit and 3-bit DQT implementation, we can observe some outliers in the blue line and red line, indicating the difficulty of training low-bit models. In contrast, the higher-bit implementations exhibit more stability throughout the training process.
4.6 Inference with Ternary Weights
Since the straight-through estimator is not employed in our proposed DQT, both the forward and backward operates directly on $n$-bits weight matrices , which means that our DQT models trained with larger bit-width are not inherently ternary during inference. To enable a fair comparison with BitNet b1.58, we adapt our models to perform ternary inference. This is achieved by using ternary weights during the forward pass while maintaining $n$-bit weights for the backward pass, updated through the straight-through estimator.
To illustrate the inference performance differences between BitNet b1.58 and our DQT models, we present their loss and perplexity scores by evaluating on the test set of WikiText-2 Merity et al. (2016)  [^4] https://huggingface.co/datasets/Salesforce/wikitext/viewer/wikitext-2-v1/test. The results are shown in Table 1. For consistency, we set the sequence length of WikiText-2 to 512, aligning with our training set. As shown in Table 1, the evaluation loss difference on WikiText-2 between our DQT 8-bit model and BitNet is minimal, with a degradation of about only 5%. When ternary inference is applied, the performance slightly decreases compared to 8-bit inference but remains close, demonstrating the robustness of our approach with ternary inference and its flexibility in deployment.
Methods Loss PPL
FP32 (reproduced) 5.39 41.93
BitNet b1.58 (reproduced) 5.52 45.83
DQT 1.58 bits (ours) 6.20 73.41
DQT 8 bits (ours) 5.80 55.75
DQT 8 bits (ternary Inf.) 5.93 60.98
Table 1: Evaluation results on WikiText-2. Loss and perplexity metrics are reported for comparison. indicates models trained with 8 bits but inference with 1.58 bits.
@@ 5 Conclusion
In this work, we explore Direct Quantized Training, a modified QAT method that directly updates low-bit weight matrices without relying on high-precision weights during training, thereby saving memory footprints. Experimental results demonstrate that DQT enables training without updating on high-precision weights, which are required for straight-through estimation. Furthermore, when extending the bit width to 8 bits, DQT models achieve performance levels comparable to FP32 models and BitNet b1.58 in both training and inference.
6 Future Work
First, we plan to extend our experiments with DQT to larger LLMs and datasets as more computational resources become available. We are particularly interested in evaluating the applicability of DQT on larger models, such as 7B parameters.
Next, the current implementation of DQT is simulated on general GPUs. For instance, we apply stochastic rounding on weight matrices optimized by the AdamW optimizer, which retains two high-precision values as momentum and variance for each element in the weight matrices. Besides, the low-bit weight matrices in our experiments are in the FP32 format, following publicly available code implementations [^5] https://huggingface.co/1bitLLM/bitnet_b1_58-large/tree/main to remain GPU-agnostic. As a result, we have not yet observed actual memory savings when comparing DQT to traditional methods like BitNet. Nevertheless, we can confirm that the weight matrices after backpropagation are indeed constrained to $n$-bits, as described in the preceding sections. Furthermore, the stored model parameters are also restricted to $n$-bits, aligning with the design of DQT. Moving forward, we aim to refine the implementation of DQT to achieve genuine memory savings and utilize simpler optimizers, e.g., SGD, to reduce the memory footprints required during optimization.
Besides, while general hardware is mainly optimized for higher-bit arithmetic operations, the support for lower bit-width operations remains sparse. This limits the practical benefits of DQT on existing hardware. In the current implementation of DQT, the low-bit computations are simulated on general GPUs that lack specialized support for ternary or other low-bit arithmetic. If GPUs were equipped with dedicated low-bit arithmetic units designed specifically for ternary or binary operations, we could expect substantial gains in both memory efficiency and computational speed. We welcome further discussion and insights from the community.

### v2

@@ 4.4 Impact of the Bit Width in DQT
Figure 4: Comparison of different bit widths in DQT. Higher $n$-bit results in better performance.
@@
Next, we proceed to examine the impact of varying the number of bits in DQT. Specifically, we experiment with $n$ in {1.58, 3, 4, 8} on 130M models trained with the Wikipedia dataset and 1B models trained with the FineWeb dataset. The corresponding results are presented in Figure 4. We select the optimal learning rate for each model.
@@
From Figure 4, we can first observe a clear trend: as the number of bits used in DQT increases, the model’s performance improves consistently for both 130M and 1B models. Notably, for the relatively lower bits DQT implementations (1.58-bit and 3-bit), we can observe some outliers in the blue line and red line, indicating the difficulty of training low-bit models. In contrast, the higher-bit implementations exhibit more stability throughout the training process.
4.5 Evaluation Results
Models Wikitext2($\downarrow$) WinoGrande ($\uparrow$) ARC (easy) ($\uparrow$) ARC (challenge) ($\uparrow$) PIQA ($\uparrow$) SciQ ($\uparrow$)
1B models (wiki)
FP32 27.03 49.01 39.56 23.29 56.20 70.70
BitNet b1.58 34.90 51.38 36.74 22.95 54.68 69.40
DQT 8 bit 30.94 49.96 37.54 23.55 56.47 69.50
DQT 8 bit (ternary Inf.) 35.51 51.30 36.36 24.23 54.19 68.90
1B models (fineweb)
FP32 19.99 52.33 48.06 25.00 67.90 80.20
BitNet b1.58 28.20 51.22 44.36 22.53 65.40 75.40
DQT 8 bit 25.43 51.97 45.12 23.63 66.38 75.90
DQT 8 bit (ternary Inf.) 27.32 51.70 45.62 22.87 65.51 75.70
Table 1: Evaluation results on different tasks. Except for WikiText-2, where perplexity is reported, we report the accuracy metric for the remaining tasks.
In this section, we conduct evaluation on the WikiText-2 test set Merity et al. (2016) and several other zero-shot tasks. We focus on 1B parameter models to best demonstrate the capability of our approach. The results are summarized in Table 1. For consistency, we set the sequence length to 512 across all tasks, matching the configuration used during training. In the table, ‘ternary Inf.’ refers to DQT models trained with 8-bit weights but evaluated using ternary weights. Details on how ternary inference is implemented for DQT models are provided in Appendix A.2. As shown in Table 1, FP32 models achieve the highest overall performance across tasks, followed closely by our DQT 8-bit variants. Notably, except for the WinoGrande task using 1B models trained on the Wikipedia dataset, our DQT 8 bit models outperform BitNet b1.58 across all other benchmarks. These results highlight that DQT 8 bit models more closely approximate the performance of FP32 models compared to BitNet. Moreover, when ternary inference is applied, the performance slightly decreases compared to 8-bit inference but remains on par with BitNet, demonstrating the robustness of our approach with ternary inference and its flexibility in deployment.
5 Analysis
While stochastic rounding offers advantages in keeping low-precision weights, it may bring some problems. For example, small weight updates may be applied with low probability (e.g., a value like -1+0.02 has only a 2% chance of being rounded to 0). If such rare updates do occur, we are particularly interested in understanding their impact on the overall training dynamics. In this section, we provide a detailed analysis of the effects and implications of using stochastic rounding.
5.1 The Role of Stochastic Rounding in DQT
Figure 5: Comparison between DQT 1.58 bits and a variant using absmax quantization for weight updates. Both methods use the same learning rate.
Firstly, we provide empirical evidence for the critical role of stochastic rounding in preserving gradient information during training with low-precision weights. Particularly, we compare 130M ternary DQT with a variant that simply uses absmax quantization on the updated weight matrices and maintains the weights in ternary format without stochastic rounding. As shown in Figure 5, the latter fails to converge, despite operating under the same bit budget. This outcome is expected, as absmax quantization can easily ignore small updates on the weights. In contrast, stochastic rounding not only performs quantization but also facilitates the accumulation of fine-grained updates, enabling effective training even in extremely low-bit regimes.
5.2 Quantifying Weight Update Frequency in DQT
Figure 6: Percentage of updated weights after each training step.
As discussed in Section 5.1, stochastic rounding facilitates training by allowing even small weight updates to take effect. To better understand its impact, we examine how frequently weights are updated during training. Specifically, we analyze 130M-size models to measure the percentage of quantized weights that change after each training step.
We quantify the weight update frequency in Figure 6 for three variants: DQT 1.58 bit, BitNet b1.58 and DQT 8 bit under the same learning rate and batch size. Note that we show the average percentages of all weight matrices in the model. For BitNet, the peak weight update rate is approximately 0.05% (observed at step 2000, the end of the warm-up phase), meaning only 0.05% of quantized ternary weights change after a single step. Ternary DQT exhibits a similar update rate of around 0.04%, indicating minimal difference between the two in terms of update frequency. In contrast, the 8-bit DQT variant, with weight values ranging from -128 to 127, shows a significantly higher update frequency, reaching up to 8%. A more detailed discussion on how we acquire these percentages is provided in Appendix A.4.
5.3 Impact of Small Weight Updates in DQT Training
Figure 7: Force to remain refers to the variant that ignores the smallest 20% of weight updates, while force to update enforces changes in this bottom 20%. All variants use the same learning rate.
Finally, to investigate the impact of small weight updates during training, we rank the absolute value of weight updates at each step and selectively intervene in the lowest 20%. For these bottom 20% updates, we apply one of two interventions: either suppress the update by retaining the original ternary value, or enforce a change by rounding it to a different quantized value, even if the update is small. For example, if an update of +0.02 from -1+0.02 falls within the smallest 20%, we either keep it at -1 (suppress) or round it to 0 (enforce), depending on the experimental condition. Figure 7 shows the training loss for these variants, where force to remain refers to suppressing the bottom 20% of updates, and force to update corresponds to enforcing them to change. From Figure 7, we observe that the original implementation of DQT (1.58-bit) achieves the best performance. Ignoring the smallest 20% of updates has minimal impact, while enforcing these small updates to take effect slightly accelerates convergence. However, the final losses at the end of training are similar across all variants. This may be due to the fact that small updates contribute less to training overall, and even when such updates do occur, stochastic rounding may re-adjust them toward the optimal value in the next training steps.
@@ 6 Conclusion
In this work, we explore Direct Quantized Training, a modified QAT method that directly updates low-bit weight matrices without relying on high-precision weights during training. This design enables DQT to operate effectively even under constrained GPU memory settings. Experimental results across different sizes of models demonstrate that DQT enables training without updating on high-precision weights, which are required for straight-through estimation. Moreover, when using 8-bit weights, DQT achieves performance comparable to both FP32 models and BitNet b1.58 during training and inference, demonstrating its effectiveness and practicality for efficient model training.
Limitations
While DQT demonstrates strong performance under low-precision and memory-constrained training environments, there are several limitations to be addressed in future work. First, although DQT achieves competitive results compared to high-precision baselines, it has not yet been tested on very large-scale models, e.g., 7B+ models. We plan to extend our approach to such scales when sufficient computational resources become available. In addition, exploring DQT’s applicability to extreme low-precision scenarios (e.g., on hardware that supports true low-bit integer calculations) is also an open direction for future work.
Appendix A Appendix
A.1 Implementation Details
Model configuration.
The detailed configurations for models of different sizes are provided in Table 2. Note that the batch size varies across model sizes during training, but no gradient accumulation is used in any experiment. For each model, the learning rate is selected via grid search over the set {1e-5, 1e-4, 5e-4, 1e-3} using our development set. Regarding the tokenizer, we adopt a publicly released pre-trained one [^5] https://huggingface.co/1bitLLM/bitnet_b1_58-large without further updates during training. We fix the random seed to 42 for reproducibility.
Dataset preprocessing.
The maximum length of training data for both datasets is set to 512. Texts longer than 512 tokens are split into separate chunks, while shorter texts are padded accordingly. For Wikipedia, this preprocessing results in approximately 14 million sentences derived from the original 6.4 million examples, while for FineWeb, the dataset yields around 33 million sentences. The test set of WikiText-2 is from  [^6] https://huggingface.co/datasets/Salesforce/wikitext/viewer/wikitext-2-v1/test.
Low-precision environments.
For FP8 simulation, we choose to use MS-AMP because the GH200 Superchips are equipped with ARM-based CPUs, which are not fully supported by the transformer-engine library currently. As for the reason for using low-precision simulation, hardware and software constraints, such as the inability to modify low-level PyTorch kernels or implement true integer-based weight updates, make it infeasible to perform actual low-bit computation in our environment. These limitations are typical in academic settings, where direct access to low-level hardware accelerators is restricted. As an alternative, we focus on evaluating the practical effectiveness of our method under memory-constrained conditions, simulating the scenarios where computational resources are limited.
Figure 8: An example of our training process using 8-bit quantization.
Params hidden_size intermediate_size num_hidden_layers num_attention_heads batch_size
130M 768 2048 12 12 64
320M 1024 2048 24 16 32
1B 2048 3072 24 32 16
Table 2: Configuration of different models sizes.
A.2 Ternary Inference
Since the straight-through estimator is not employed in our proposed DQT, both the forward and backward operate directly on $n$-bit weight matrices, which means that our DQT models trained with larger bit-width are not inherently ternary during inference. To enable a fair comparison with BitNet b1.58, we adapt our models to perform ternary inference. This is achieved by using ternary weights during the forward pass while maintaining $n$-bit weights for the backward pass, updated through the straight-through estimator. The straight-through estimator is employed only to enable ternary inference in $n$-bit DQT models as a variant of our models. We additionally present the training loss for DQT 8 bit and DQT 8 bit with ternary inference in Figure 9.
Figure 9: Training loss comparison of DQT 8 bits and DQT 8 bits that utilizes ternary inference. DQT achieves ternary inference with minimal degradation.
Figure 10: Non-logarithmic training loss comparison of DQT 8 bits and BitNet b1.58 in 1B sizes. Our DQT 8 bits performs slightly better than BitNet b1.58.
A.3 Actual Memory Usage
We present the actual memory usage of models with different sizes in Table 3. The values reflect usage on a single GH200 GPU, which has 97,871 MB of available memory. For 130M models, we use 4 GPUs, and for 1B models, 16 GPUs are used. We can also observe exactly how much memory is saved after applying Adafactor in BF16 and FP8 for different sizes of models from Table 3.
A.4 Weight Update Frequency
For BitNet models, we compare quantized weight matrices at adjacent training steps by iterating through all corresponding rows and columns to identify whether each element is updated or not. For DQT models, we can directly compare the weight matrices before and after stochastic rounding. The update percentage is computed as the ratio of changed weight elements to the total number of elements.
Model size FP32 BF16 BF16+Adafactor FP8 FP8+Adafactor
130M 69327 54675 53827 39276 38315
1B 76533 58345 53723 40945 37669
Table 3: Actual GPU memory usage (in MB) of models with different sizes on a single GH200.